40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Mon Dec 13 18:50:02 2010, frank, tara, DailyProgress, Foam, increase inner insulation holes' size 100_1176.jpg100_1179.jpg
    Reply  Tue Dec 14 22:00:01 2010, frank, tara, DailyProgress, Foam, increase inner insulation holes' size beat_2010_12_14.png
Message ID: 429     Entry time: Tue Dec 14 22:00:01 2010     In reply to: 428
Author: frank, tara 
Type: DailyProgress 
Category: Foam 
Subject: increase inner insulation holes' size 

After we fixed the insulation, the noise spectrum at low frequency became better. 

 

I plotted the result with LO noise. The LO phase noise is now divided by sqrt of 2 for beating two LOs together.

* the legend in the plot is wrong, it should be for "100kHz input range" not 10kHz

 

 The measurement was done between IFR2023A and B, they probably have

different noise level. so the curves do not exactly match.

 

For the red plot, the data was taken before I cleaned the lens behind the 2nd faraday isolator, and re-alinged the beam to RCAV.

After that I took the data for blue and green plots. I'm not sure why they are not the same, probably because Frank and Peter were walking around the lab. 

 

Peaks at 3.75Hz, 8 Hz  are cavities' pendulum resonance frequency. When I push the table, the peaks get higher.

I think dirty optics might still cause extra noise. I'll try to clean ones that are still dirty, and see if there will be any improvement.

Quote:

Today we increased the size of the beam path on inner insulations for ACAV and RCAV, cleaned vacuum chambers' windows, and optics on the table.

 

We have suspected that scattering light dominates the noise  spectrum at low frequency, so we checked all possible scattering sources from the vacuum chamber.

 

We increased the size of the opening for the beam paths on ACAV/RCAV insulation caps.

We removed the cap by cutting the AL tape around, after that the caps are just placed back. We haven't taped them yet.

 ACAV's chamber windows are especially dirty, full of scratches, so we cleaned them with methanol.

We also found a reflection coming back from ACAV. The spot size is quite large, ~order of cm. The position is quite far from the incoming beam.

The source is still undetermined.

 

Optics on the setup are cleaned. They had dust, finger prints on them.

 

ACAV's periscope was readjusted, so that the beam hits near the center of the mirrors, not close to the edge of the mirror.

 

 RCAV temperature sensors are fixed and not noisy anymore.

Plots and pictures on the board outside the lab are updated.

 

As we are waiting for temperature to settle, we clean the lab.

 

Attachment 1: beat_2010_12_14.png  41 kB  | Hide | Hide all
beat_2010_12_14.png
ELOG V3.1.3-